Please help me solve and check this equation. Thank you!
1/2x=22

Answer by checkley79(3341)

1/2x=22
X/2=22 CROSS MULTIPLY.
X=2*22
X=44 ANS.
PROOF:
1/2*44=22
44/2=22
22=22